OpenStudy (kirbykirby):

If f(x) = integral [cos(t^2)dt] from 0 to x, using the known power series for cos x, find a power series for f(1). This is f(x): \[f(x)=\int\limits\limits_{0}^{X}\cos(t^2)dt\]

OpenStudy (kirbykirby):

I ended up finding \[\sum_{n=0}^{infinity}\frac{(-1)^{n}x^{4n+1}}{(4n+1)(2n)!}=x-\frac{x^5}{5*2!}+\frac{x^9}{9*4!}-\frac{x^{13}}{13*6!}+...\] But I'm not quite sure what it means to find the power series for f(1).. Isn't that just a number (whatever I found evaluated at 1)? There is no more x-dependence so how is it a power series :S I'm so confused

OpenStudy (anonymous):

maybe its a power series centered at x = 1

OpenStudy (kirbykirby):

Hmm maybe that's what they meant..

OpenStudy (anonymous):

You just plug in \(x-1\). Centering is not an issue here, because we're dealing with Maclaurin series which are always centered at \(0\).

OpenStudy (anonymous):

I mean \(x=1\)